# DiagGPT: An LLM-based and Multi-agent Dialogue System with Automatic Topic Management for Flexible Task-Oriented Dialogue

### v3

@@ Abstract
Large Language Models (LLMs), such as ChatGPT, are increasingly sophisticated and exhibit capabilities closely resembling those of humans. A significant application of these LLMs is their use as chat agents, responding to human inquiries across various domains. While current LLMs proficiently answer general questions, they often fall short in complex diagnostic scenarios such as legal, medical, or other specialized consultations. These scenarios typically require Task-Oriented Dialogue (TOD), where an AI chat agent must proactively pose questions and guide users toward specific goals or task completion. Previous fine-tuning models have underperformed in TOD and the full potential of this capability in current LLMs has not yet been fully explored. In this paper, we introduce DiagGPT (Dialogue in Diagnosis GPT), an innovative approach that extends LLMs to more TOD scenarios. In addition to guiding users to complete tasks, DiagGPT can effectively manage the status of all topics throughout the dialogue development. This feature enhances user experience and offers a more flexible interaction in TOD. Our experiments demonstrate that DiagGPT exhibits outstanding performance in conducting TOD with users, showing its potential for practical applications in various fields.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT, have demonstrated remarkable performance on various natural language processing (NLP) tasks Brown et al. (2020); Chowdhery et al. (2022); Wei et al. (2022a); OpenAI (2023). In many cases, OpenAI GPT-4 even outperforms human performance OpenAI (2023). With the use of prompt engineering techniques (e.g., chain-of-thought prompting Brown et al. (2020); Wei et al. (2022b), in-context learning Brown et al. (2020); Xie et al. (2022); Min et al. (2022), etc.), we can unlock the unlimited potential of LLMs to complete complex tasks in our daily life. LLMs have attracted enormous attention from both academia and industry, inspiring more people to build fantastic applications based on them.
@@
However, dialogue scenarios in our daily life can be more complex. For instance, in specialized professional consultation scenarios like legal or medical diagnosis, the AI model needs to consider the user’s unique situation or information. In the process of obtaining user information, the interactive experience provided by the AI model is also crucial. The system needs to proactively ask questions. Therefore, we need a consultation process from the AI model that better simulates real medical experts and legal professionals. The AI model should conduct question-answering, topic management, and guiding users towards specific goals or task completion. This type of dialogue is known as Task-Oriented Dialogue (TOD). There are usually some predefined goals in a conversation. TOD helps users achieve their specific goals, focusing on understanding users, tracking states, and generating next actions Balaraman et al. (2021). It is substantially different from light-conversational or open-domain dialogue scenarios. Despite much research in this area, it remains challenging due to issues such as a lack of training data, inefficiency, and drawbacks of fine-tuning small models, including an inability to fully understand user meaning and poor generative performance. Models from the existing research on this topic is not robust and universal. For example, fine-tuning models require a lot of data to train and are difficult to transfer to other scenarios. On the other side, although LLMs have a wide range of knowledge and the quality of their answer is far beyond that of fine-tuning models, traditional LLMs no longer meet needs of TOD and cannot effectively manage complex dialogue logic. Because they maintain a simple memory base and can only handle linear interaction.
@@
Recent advancements have focused on using LLMs as AI agents to construct multi-agent systems or to teach AI how to use tools to accomplish more complex tasks Schick et al. (2023); Shen et al. (2023). These systems typically have a core AI agent that controls the entire task process. A prominent example is AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, which employs multiple GPT models to strategize the responsibilities of each agent in order to split complex tasks and then complete them. In such multi-agent systems, the key lies in the division of tasks and the interaction between agents. By organizing multiple LLMs and instruct them to collaborate, we have the opportunity to tackle many complex tasks which one LLM cannot do well.
@@
Fine-tuning models fall short in terms of scenario transfer ability and training data requirements, while a single large language model is not proficient in dialogue state tracking and management. However, we can leverage the strong knowledge background of LLMs and employ a multi-agent framework to incorporate the dialogue state tracking and management ideas from fine-tuning models. Thus, we can construct a multi-agent dialogue system that fulfills the requirements of task-oriented dialogue.
@@
  • Question Answering: It is the distinctive feature of traditional LLM-based conversational systems. LLMs possess a wide range of knowledge and can provide high-quality answers to various questions. In our system, we retain this fundamental ability of LLMs.
  • Task Guidance: The system is designed to guide users towards a specific goal and assist them in accomplishing the task throughout the dialogue progression. This is achieved by advancing a sequence of predefined topics throughout the dialogue.
  • Proactive Asking: The system has the ability to proactively pose questions based on a predefined checklist, thereby collecting necessary information from users.
  • Topic Management: The system is capable of automatically managing topics throughout the dialogue, tracking topic progression, and effectively engaging in discussions centered around the current topic. It performs well in managing various topic changes in complex dialogues.
  • Versatile: Our system is directly based on LLMs. It can perform well in various scenarios without requiring any training data, a capability that previous fine-tuning models lack. This system can be easily applied to multiple cases by defining specific predefined goals and supplementing functions to support them.
  • High Extendibility: In this paper, we only introduce the basic framework of this AI system aimed at achieving task-oriented dialogue. We have designed the system with ample flexibility to incorporate additional functions to handle tasks in complex scenarios and to meet more needs of conversational systems. Besides, as the foundation model is upgraded, the performance of our system will also become better.
Given these features, DiagGPT can meet the aforementioned needs and better engage in professional consultation conversations with users in complex scenarios. Our main contribution is to make traditional LLM-based conversational systems smarter in flexible task-oriented dialogue. We build on the strong knowledge of LLMs and give them more interactive capabilities. Therefore, DiagGPT can function like a more intelligent and more professional chatbot.  [^3] Our code will be made publicly available after the paper’s acceptance.
2 Related Works
@@ 3.1 DiagGPT Framework
DiagGPT is a multi-agent and collaborative system composed of several modules: Chat Agent, Topic Manager, Topic Enricher, and Context Manager. Each module is a LLM with specific prompts that guide their function and responsibility. Among these modules, the Topic Manager is particularly important as it tracks the dialogue state and automatically manages the dialogue topic.
@@ 3.3 Maintaining Topic Stack
The topic stack is a data structure in this AI system that stores and tracks the dialogue state. We consider the progress of a dialogue to have multiple stages or states, and these states follow a first-in, first-out (FIFO) order, which can be effectively modeled using a stack. Although we refer to this topic storage structure as a stack, this component does not strictly adhere to the FIFO rule. This is due to the presence of complex dialogue logic, and FIFO is merely the most common form of topic development. We will provide a detailed description of some operations of this structure in the following paragraphs.
@@
Figure 3: The action of creating a new topic.
Figure 4: The action of finishing the current topic.
@@
Figure 5: The action of jumping to an existing topic.
Round Count Completion Rate Success Rate Dialogue Quality
ChatGPT (gpt-3.5-turbo) 10.0 0.8 0.6 6.8
ChatGPT (gpt-4) 5.3 1.0 1.0 9.1
DiagGPT (Ours) 5.6 1.0 1.0 9.0
Table 1: Quantitative experimental results. We compare our methods DiagGPT with ChatGPT (gpt-turbo-3.5 and gpt-4) in 20 different kinds of TOD scenarios. The results show that DiagGPT can perform the traditional TOD well.
@@ 3.4 Enriching Topic
We select the top item in the topic stack as the current topic. However, it cannot be directly used as a chat topic for the Chat Agent to interact with the user. The Topic Enricher is designed to bridge this gap and assist in better organizing the language for use. We initially categorize the topic into Ask user and Answer user. Typically, newly generated topics fall under Answer user, while predefined topics are categorized as Ask user. This distinction helps the system determine whether to answer a user’s question or ask a question in the current round of dialogue. The Topic Enricher takes the output of the Context Manager and the current topic to enrich it into a topic that contains ample information and is contextually appropriate. This enriched topic is then provided to the Chat Agent.
@@ 4 Experiments
We conduct both quantitative experiments and qualitative experiments to demonstrate the performance of DiagGPT. Details of the system implementation can be found in Appendix C.
The quantitative evaluation of DiagGPT is challenging due to the difficulty in quantifying the system’s ability to manage various topics that develop in a dialogue. Therefore, we employ quantitative experiments to only assess the system’s basic capabilities in topic advancement and guidance. Additionally, we do not evaluate our system using task-oriented dialogue datasets such as DialoGLUEYadav et al. (2019); Moghe et al. (2023), as DiagGPT is an open system that relies heavily on AI consciousness combined with prompt instructions for output. Unlike fine-tuning models, our system does not require training. We have developed a new dataset called LLM-TOD, which encompasses 20 different topics in task-oriented dialogue scenarios (details are shown in Appendix F). Each dataset entry includes a checklist of items to go through, finally achieving a specific goal. This checklist and the goal are input into the LLMs to complete a TOD. A LLM-based user simulator is also employed to interact with the dialogue system. The results are evaluated using a combination of human assessment and another LLM.
On the other hand, qualitative experiments more effectively demonstrate the capabilities and performance of DiagGPT. As shown in Appendix D, these experiments show that DiagGPT can conduct task-oriented dialogues more flexibly.
@@ 4.1 Quantitative Experiment Results
From Table 1, we observe the results of the quantitative experiment. We evaluated the performance from four perspectives: average rounds needed to complete a dialogue (Round Count), the percentage of completed checklist items out of all items (Completion Rate), whether the system successfully completed the task to achieve the goal (Success Rate), and Dialogue Quality (0-10), which is a average scored and is assessed using an LLM to evaluate the overall chat history. This table suggests that DiagGPT can match the performance of the state-of-the-art LLM, gpt-4, and effectively perform traditional TOD. Our advantage compared to gpt-4 is that DiagGPT interacts with users flexibly through its automatic topic management. This feature is evident in the qualitative experiment.
@@ 5 Conclusion
In this paper, we propose DiagGPT, an LLM-based and multi-agent system designed to complete task-oriented dialogue tasks. The principle of our system is to leverage the strong understanding and reasoning capabilities of Large Language Models to design an AI agent that can automatically manage topics and track dialogue state. This feature enables our system to accurately comprehend users’ intentions, assisting them in completing specific tasks more flexibly. The system is particularly well-suited for real-world consulting scenarios in flexible TOD. DiagGPT demonstrates the unlimited potential of LLMs in more complex dialogue scenarios, such as task-oriented dialogues, benefiting society by applying AI in various scenarios.
In the future, we aim to explore how to better serve users by combining the robust capabilities of LLMs. Moreover, we believe that the construction of LLM-based multi-agent systems signifies the future of AI development. We hope that the design of our system can inspire the development of more sophisticated AI applications and pave the way for LLMs towards more advanced AI systems.
Appendix A Limitation
@@
Appendix B Analysis of DiagGPT Extendibility

### v4

@@ Abstract
A significant application of Large Language Models (LLMs), like ChatGPT, is their deployment as chat agents, which respond to human inquiries across a variety of domains. While current LLMs proficiently answer general questions, they often fall short in complex diagnostic scenarios such as legal, medical, or other specialized consultations. These scenarios typically require Task-Oriented Dialogue (TOD), where an AI chat agent must proactively pose questions and guide users toward specific goals or task completion. Previous fine-tuning models have underperformed in TOD and the full potential of conversational capability in current LLMs has not yet been fully explored. In this paper, we introduce DiagGPT (Dialogue in Diagnosis GPT), an innovative approach that extends LLMs to more TOD scenarios. In addition to guiding users to complete tasks, DiagGPT can effectively manage the status of all topics throughout the dialogue development. This feature enhances user experience and offers a more flexible interaction in TOD. Our experiments demonstrate that DiagGPT exhibits outstanding performance in conducting TOD with users, showing its potential for practical applications in various fields.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT, have demonstrated remarkable performance on various natural language processing (NLP) tasks Brown et al. (2020); Chowdhery et al. (2022); Wei et al. (2022a); OpenAI (2023). In many cases, OpenAI GPT-4 even outperforms human performance OpenAI (2023). With the use of prompt engineering techniques (e.g., chain-of-thought prompting Brown et al. (2020); Wei et al. (2022b), in-context learning Brown et al. (2020); Xie et al. (2022); Min et al. (2022), etc.), we can improve reasoning and understanding of LLMs to complete complex tasks in our daily life. LLMs have attracted enormous attention from both academia and industry, inspiring more people to build useful applications based on them.
@@
However, dialogue scenarios in our daily life can be more complex. For instance, in specialized professional consultation scenarios like legal or medical diagnosis, the chat agent needs to consider the user’s unique situation or information. In the process of obtaining user information, the interactive experience provided by the agent is also crucial. The system needs to proactively ask questions. Therefore, we need a consultation process from the chat agent that better simulates real medical experts and legal professionals. The chat agent should conduct question-answering, topic management, and guiding users towards specific goals or task completion. This type of dialogue is known as Task-Oriented Dialogue (TOD). There are usually some predefined goals in a conversation. TOD helps users achieve their specific goals, focusing on understanding users, tracking states, and generating next actions Balaraman et al. (2021). It is substantially different from light-conversational or open-domain dialogue scenarios. Despite much research in this area, it remains challenging due to issues such as a lack of training data, inefficiency, and drawbacks of fine-tuning small models, including an inability to fully understand user meaning and poor generative performance. Models from the existing research on this topic is not robust and universal. For example, fine-tuning models require a lot of data to train and are difficult to transfer to other scenarios. On the other side, although LLMs have a wide range of knowledge and the quality of their answer is far beyond that of fine-tuning models, traditional LLMs no longer meet needs of TOD and cannot effectively manage complex dialogue logic. Because they maintain a simple memory base and can only handle linear interaction.
@@
Recent advancements have focused on using LLMs as agents to construct multi-agent systems or to teach LLMs how to use tools to accomplish more complex tasks Schick et al. (2023); Shen et al. (2023). These systems typically have a core agent that controls the entire task process. A prominent example is AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, which employs multiple GPT models to strategize the responsibilities of each agent in order to split complex tasks and then complete them. In such multi-agent systems, the key lies in the division of tasks and the interaction between agents. By organizing multiple LLMs and instruct them to collaborate, we have the opportunity to tackle many complex tasks which one LLM cannot do well.
@@
Fine-tuning models fall short in terms of scenario transfer ability and massive training data requirements, while a single large language model is not proficient in dialogue state tracking and management. However, we can leverage the strong knowledge background of LLMs and employ a multi-agent framework to incorporate the dialogue state tracking and management ideas from fine-tuning models. Thus, we can construct a multi-agent dialogue system that fulfills the requirements of flexible task-oriented dialogue.
@@
  • Question Answering It is a basic feature of traditional LLM-based conversational systems. LLMs possess a wide range of knowledge and can provide high-quality answers to various questions. we build upon and retain this essential capability of LLMs.
  • Task Guidance The system is designed to guide users towards a specific goal and assist them in accomplishing the task throughout the dialogue progression. This is achieved by advancing a sequence of predefined topics throughout the dialogue.
  • Proactive Asking The system has the ability to proactively pose questions based on a predefined checklist, thereby collecting necessary information from users.
  • Topic Management The system is capable of automatically managing topics throughout the dialogue, tracking topic progression, and effectively engaging in discussions centered around the current topic. It performs well in managing various topic changes in complex dialogues.
  • Versatile Our system is directly based on LLMs. It can perform well in various scenarios without requiring any training data, a capability that previous fine-tuning models lack. This system can be easily applied to multiple cases by defining specific predefined goals and supplementing functions to support them.
Given these features, DiagGPT can meet the aforementioned needs and better engage in professional consultation conversations with users in complex scenarios. Our main contribution is to make traditional LLM-based conversational systems smarter in flexible task-oriented dialogue. We build on the strong knowledge of LLMs and give them more interactive capabilities. Therefore, DiagGPT can function like a more intelligent and more professional chatbot.
2 Related Work
@@ 3.1 DiagGPT Framework
DiagGPT is a multi-agent and collaborative system composed of several modules: Chat Agent, Topic Manager, Topic Enricher, and Context Manager. Each module is a LLM with specific prompts that guide their function and responsibility. Among these modules, the Topic Manager is the most important as it tracks the dialogue state and automatically manages the dialogue topic.
@@ 3.3 Maintaining Topic Stack
The topic stack is a data structure in this system that stores and tracks the dialogue state. Stack management is a kind of simulation of dialog state management in both short-term and long-term context. In our system, we concretize the dialogue state of the LLM and explicitly store it, allowing the system to perform better over the long term context and avoid ambiguity and forgetfulness. We consider the progress of a dialogue to have multiple stages or states, and these states follow a first-in, first-out (FIFO) order, which can be effectively modeled using a stack. Although we refer to this topic storage structure as a stack, this component does not strictly adhere to the FIFO rule. This is due to the presence of complex dialogue logic, and FIFO is merely the most common form of topic development. We will provide a detailed description of some operations of this structure in the following paragraphs.
@@
Figure 3: The action of create a new topic.
Figure 4: The action of finish the current topic.
@@
Figure 5: The action of jump to an existing topic.
Human Evaluation LLM Evaluation
RC $\downarrow$ CR $\uparrow$ SR $\uparrow$ CS $\uparrow$ CR $\uparrow$ SR $\uparrow$ RQ $\uparrow$ CS $\uparrow$
ChatGPT (gpt-3.5-turbo) 9.0 0.8 0.4 - 0.8 0.7 7.8 -
ChatGPT (gpt-4) 7.7 1.0 1.0 5.0 1.0 1.0 9.0 8.5
DiagGPT (Ours) 7.0 1.0 1.0 15.0 1.0 1.0 9.0 11.5
Table 1: Quantitative experimental results. We compare our method DiagGPT with ChatGPT (gpt-turbo-3.5 and gpt-4) across 20 diverse dialogue scenarios, assessed by Round Count (RC), Completion Rate (CR), Response Quality (RQ), and Comparison Score (CS).
@@ 3.4 Enriching Topic
We select the top item in the topic stack as the current topic. However, it cannot be directly used as a chat topic for the Chat Agent to interact with the user. Topics in the topic stack are simple and convenient to store, whereas topics for the practical use of the Chat Agent need to include more information. Without a topic enricher, it is difficult for the Chat Agent to directly understand the main objective in the current round of dialogue.
The Topic Enricher is designed to bridge this gap and assist in better organizing the language for use. We initially categorize the topic into Ask user and Answer user. Typically, newly generated topics fall under Answer user, while predefined topics are categorized as Ask user. This design helps the system determine whether to answer a user’s question or ask a question in the current round of dialogue. The Topic Enricher takes the output of the Context Manager and the current topic to enrich it into a topic that contains more context information. This enriched topic is then provided to the Chat Agent.
@@ 4 Experiments
4.1 Setups
We conduct both quantitative experiments and qualitative experiments to demonstrate the performance of DiagGPT. Details of the system implementation can be found in Appendix B.
The quantitative evaluation of DiagGPT is challenging due to the difficulty in quantifying the system’s ability to manage various topics that develop in a dialogue. Therefore, we employ quantitative experiments to only assess the system’s basic capabilities in topic advancement and guidance. We do not evaluate our system using task-oriented dialogue datasets such as DialoGLUE Yadav et al. (2019); Moghe et al. (2023). As DiagGPT is an open system, and it does not require training unlike fine-tuning models. The output of our system is conditioned solely on prompts. Moreover, these traditional TOD datasets focuses on specific subtasks (intent prediction, slot filling, etc.) that were considered necessary for TOD, while LLM-based systems circumvents these specific tasks in a more end-to-end fashion. Therefore, evaluating our system using traditional methods is challenging.
To quantitatively evaluate our system, we develop a new dataset called LLM-TOD, which encompasses 20 different topics in task-oriented dialogue scenarios (details are shown in Appendix C). Each data of a TOD scenario includes a checklist of items to go through, finally achieving a specific goal. This checklist and the goal are input into the LLMs to complete a TOD. Besides, we develop a new evaluation schema based on LLMs for TOD tasks. An LLM-based user simulator is also employed to interact with the dialogue system. The results are evaluated using a combination of human assessment and another LLM. During the LLM evaluation, the inputs include basic TOD information and the complete chat history.
On the other hand, qualitative experiments more effectively demonstrate the capabilities and performance of DiagGPT. These experiments show that DiagGPT can conduct task-oriented dialogues more flexibly.
@@ 4.2 Quantitative Experiment Results
We evaluate the performance through both human evaluation and LLM evaluation. We compared our method with two ChatGPT baselines: gpt-3.5-turbo and gpt-4. The evaluation involved the following metrics:
  • Round Count (RC): The average number of rounds required to complete a task-oriented dialogue. This metric can be seen as a measure of system efficiency, where lower values are preferable.
  • Completion Rate (CR): The percentage of completed checklist items out of the total number of items. An ideal system should complete all items on the checklist.
  • Success Rate (SR): The indicator of whether the system successfully completed the task to achieve the intended goal. This metric is a value of 0 or 1 for each task-oriented dialogue.
  • Response Quality (RQ): A score assessed using an LLM to evaluate the quality of the entire chat history and system response. The values range from 1 to 10, as generated by an LLM.
  • Comparison Score (CS): The score of comparison between responses from ChatGPT (gpt-4) and DiagGPT within a single dialogue. It is calculated based on the total number of wins in various scenarios, indicating which model performed better.
In detail, we calculated the average score across 20 different TOD scenarios. For the comparison score from the LLM evaluation, we alternated the position of the chat history from two systems in the prompt and evaluated twice to obtain an average score. This approach helps to avoid positional bias in automatic evaluation with LLMs. For the evaluation of response quality, we consider factors such as understanding, relevance, experience, and sufficiency, etc.
From Table 1, we can observe the results of the quantitative experiment. This table shows that DiagGPT can match the performance of the state-of-the-art LLM, gpt-4, and effectively manage task-oriented dialogue in terms of response quality. Our system completes all tasks on the checklist and successfully achieves the goals in all TOD scenarios with a lower round count compared to ChatGPT. It demonstrates superior efficiency in conducting TOD by leading users to their goals and completing tasks effectively, which is crucial in TOD. Given the comparable response quality, we directly compared responses from the two systems. Both human and LLM evaluation results suggest that our system is more effective and perform better.
Our advantage over gpt-4 is that DiagGPT interacts with users more flexibly through its automatic topic management. This feature becomes more apparent in the subsequent qualitative experiment.
Figure 6: A dialogue example in the medical consulting scenario. The system DiagGPT acts as a real doctor to proactively ask for information and answer the user’s questions.
Figure 8: A dialogue example in the medical consulting scenario when the user ask some questions about COVID-19.
4.3 Qualitative Experiment Results
We modified the settings in DiagGPT to adapt it for use in medical dialogue scenarios for conducting qualitative experiments. Figure 6 and Figure 7 present a complete dialogue demonstration in the medical consulting process. This is a medical diagnosis where the task is to help the patient identify the cause and give advice. Medical consulting is not a pure open-end question-answering. Because patients usually lack extensive medical knowledge, they rely on doctors to instruct them to give their personal information. Therefore, users can experience a real doctor, not a dull medical question-answering machine. The user acts as a patient, while the system is a doctor, initially collecting information and gradually providing advice to the patient. The main dialogue development follows the checklist: Basic information, Chief complaint, Duration of symptoms, Severity of symptoms, which are also predefined topics.
This demonstration has showcased the robust conversational ability of the DiagGPT, which can proactively ask questions and guide the user to the final goal of the task, thereby achieving task-oriented dialogue. It simulates many real consulting scenarios. Other conversational systems, such as general ChatGPT, cannot achieve this performance. They usually can only answer users’ questions and find it challenging to complete specific goals in complicated scenarios, even with elaborated prompts.
4.4 Case Study of Automatic Topic Management
The core capability of DiagGPT is to automatically manage topics throughout the dialogue, which is the function and primary responsibility of Topic Manager. Figure 6 and Figure 7 illustrate the primary checklist progression. The topics from the checklist are retrieved and discussed sequentially, demonstrating the action of finish the current topic. When the conversation reaches the severity of symptoms, we observe that the dialogue topic remains here in several rounds of dialogue, allowing the system to have time on understanding the user’s conditions. This is a typical situation in real-world scenarios where users do not provide enough information for the doctor. The action of create a new topic is shown in Figure 8. Here, the user actively consults the system with some information about COVID-19 to check symptoms. We observe that the system generates a new topic about COVID-19 and discusses this, rather than rigidly following the checklist. These results all demonstrate the effectiveness of Topic Manager. These case studies fully showcase the system’s flexible understanding capabilities, demonstrating its proficient handling of various scenarios that may occur in real-world dialogues.
@@ 5 Conclusion
In this paper, we propose DiagGPT, an LLM-based and multi-agent system designed to complete task-oriented dialogue tasks in a flexible way. The principle of our system is to leverage the strong understanding and reasoning capabilities of Large Language Models to implement agents that can automatically manage topics and track dialogue state. This feature enables our system to accurately comprehend intentions of users in different situations, assisting them in completing specific tasks more flexibly. The system is particularly well-suited for real-world consulting scenarios in flexible TOD. Both quantitative and qualitative experiments demonstrate the superior performance of DiagGPT.
However, DiagGPT represents a experimental system designed to demonstrate the potential of LLMs in handling real tasks. In the future, we aim to explore how to better serve users by leveraging the robust capabilities of LLMs to achieve more functionality and complete more complicated tasks.
Limitation
@@
Appendix A Analysis of DiagGPT Extendibility